OpenStudy (anonymous):

Find the laplace transform of the function: f(t)=2 if 0 <= t < 3; f(t)=0 if t >= 3.

OpenStudy (anonymous):

u know that\[L(f)=\int_0^\infty f(t) e^{-st} dt\]and u can write for this particular f\[L(f)=\int_0^3 2 \times e^{-st} dt+\int_3^\infty 0 \times e^{-st} dt=2\int_0^3 e^{-st} dt\]

OpenStudy (anonymous):

last integral is easy to calculate :)

OpenStudy (primeralph):

follow @mukushla 's explanation

OpenStudy (anonymous):

Oh wow guys, thanks.

OpenStudy (anonymous):

yw :)